HeartTricker:

Select the exclusion that fits best with this problem. A. x = 0 B. x, y, z = 0 C. 1 + y + y 2, y = 0, 1 - y = 0 D. a, b = 0, a 2 + ab + b2 = 0 E. 2x - 1 = 0, 4x 2 + 2x + 1 = 0

HeartTricker:

1 attachment
HeartTricker:

Im leaning towards C maybe?

ItzAgar:

Which choice would you want to eliminate?

HeartTricker:

𝗔 for sure and maybe 𝗕

AZ:

The easiest way to solve this question is look at the question and then look at the answer choices. What is the variable in your question? It's 'y' That means your answer choice has to have y. It cannot have a, b, c It cannot have x either It has to have y only

HeartTricker:

Ok, so its was 𝗖. 𝟭 + 𝘆 + 𝘆 𝟮, 𝘆 = 𝟬, 𝟭 - 𝘆 = 𝟬

AZ:

When we're talking about exclusion that means we're trying to find values that are excluded from the equation And that's found by setting the denominator equal to 0 2y - 2y^4 = 0 and if you factor this out, you'll realize how you get to the correct answer

HeartTricker:

Thank you so much!!

AZ:

@hearttricker wrote:
Ok, so its was 𝗖. 𝟭 + 𝘆 + 𝘆 𝟮, 𝘆 = 𝟬, 𝟭 - 𝘆 = 𝟬
Yup

AZ:

@az wrote:
When we're talking about exclusion that means we're trying to find values that are excluded from the equation And that's found by setting the denominator equal to 0 2y - 2y^4 = 0 and if you factor this out, you'll realize how you get to the correct answer
If you factor it, you get -2(y - 1)(y)(y^2 + y + 1) = 0 and so if you set each term equal to 0, that's how you get your answer y - 1 = 0 y = 0 y^2 + y + 1 = 0